符号・格子・頂点作用素代数と有限群

研究集会

京都大学数理解析研究所の共同利用の一つとして，下記のように研究集会を催しますので，ご案内申し上げます．

研究代表者北詰正顕（千葉大・理）

記

日 時：２００１年２月１９日 (月) １０：００〜２月２１日 (水) １６：４０

場 所：京都大学数理解析研究所４階４２０号室 {京都市左京区北白川追分町 \atop 市バス農学部前または北白川下車}

プログラム

2 月 19 日 (月) 10:00–10:50 坂内英一 (九大・数理)

Modular invariants of modular data of finite groups

(坂内悦子, 島袋修, 田上真との共同研究）

11:00–11:50 吉荒聡 (大阪教育大)

A generalized quadrangle with an automorphism group acting regularly on its points

13:30–14:20 平坂貢 (浦項 (Pohang) 工科大)

Sylow’s Theorem for Association Schemes (M. Muzychuk, P.-H. Zieschang との共同研究)

14:30–15:00 城本啓介 (熊本大・学振研究員)

Higher weights for codes over finite rings

15:10–15:40 原田昌晃 (山形大・理)

Type II codes over Z_{2k} and the Leech lattice

15:50–16:20 別宮耕一 (名大・多元数理)

Classification of Type II code over GF(4) of some small lengths 16:30–17:00 田辺顕一朗 (筑波大・数学)

小関多項式と符号の被覆半径の計算について

2 月 20 日 (火) 10:00–10:50 松澤淳一 (京大・工)

3 次曲面族の幾何とルート系

11:00–11:50 杉山健一 (千葉大・理)

有限素体上定義された Modular Codes について

13:30–14:20 安部利之 (阪大・理)

頂点作用素代数 V_{L}^{+} の有理性について

14:30–15:20 松尾厚 (東大・数理)

Ｗ代数とモンスター

15:30–16:00 島倉裕樹 (東大・数理)

Moonshine module and codes over Z/6Z

16:10–16:40 山内博 (筑波大・数学)

オービフォルドモデルのモジュラー不変性

2 月 21 日 (水) 10:00–10:30 和田倶幸 (東京農工大)

Some inequalities between invariants of blocks

(B. Kuelshammer との共同研究)

10:40–11:10 石川賢太 (千葉大・自然科学)

On finite p-groups with only two conjugacy lengths

11:20–11:50 宗政昭弘 (九大・数理)

Spherical 5-designs obtained from the unitary group U_{2m}(2)

13:30–14:20 田中太初 (九大・数理)

Some Results on the Multiplicity-Free Permutation Group GL(4, q) Acting on GL(4, q)/GL(2, q^{2})

14:30–15:00 小関道夫 (山形大・理)

Combinatorial deformations of theta series of even unimodular lattices

15:10–16:00 原田耕一郎 (オハイオ州立大)

My talk V

16:10–16:40 TBA